# Ali Dasdan KD Consulting Saratoga, CA, USA

### v2

@@ 5 The Tragedy of AI Governance
Scope of Analysis: Our analysis focuses on the governance challenge posed by transformative AI breakthroughs—systems approaching or exceeding human-level general intelligence (AGI) or superintelligence. While current AI competition exhibits different dynamics (multiple viable competitors survive capability gaps, as evidenced by Google’s recovery from Transformer lag and ongoing competition after ChatGPT and DeepSeek breakthroughs), three factors justify focusing on transformative AI: First, such breakthroughs represent the coordination challenge that matters most given potential existential consequences. Second, timing remains inherently unpredictable—just as few anticipated the deep learning revolution, Transformer architecture’s impact, ChatGPT’s capabilities, or DeepSeek’s efficiency breakthrough. Third, once such systems emerge, the competitive dynamics we analyze become immediately relevant, making ex-ante coordination essential rather than reactive governance possible.
@@ 5.1.3 Condition 3: Dominance Property
The dominance is driven by existential stakes for firms, nations, and researchers. For example, a firm who cannot raise capital for massive compute (or acceleration) risk obsolescence. This structural pressure forced even the leading non-profit actor to restructure first into a “capped-profit” entity (OpenAI, 2019) and eventually into a Public Benefit Corporation to remove fundraising limits (OpenAI, 2025).
@@ 5.1.4 Condition 4: Pareto-Inefficiency
Asymmetric Gains and the Winner-Take-All Problem AI development creates potential for extremely asymmetric outcomes. Unlike productivity competition where firms maintain roughly proportional positions, AI capabilities races may produce winner-take-all dynamics where the first to achieve advanced AI captures decisive advantages, potentially controlling entire markets, achieving strategic dominance, or even technological singularity.
@@
First, existential risk is non-excludable. The catastrophic risks from misaligned or insufficiently robust advanced AI systems affect all actors regardless of who develops them first (Bostrom, 2014). A winner who achieves advanced AI through hasty development faces the same risks of loss of control, unintended consequences, or catastrophic failure as any other actor. The instrumental convergence thesis (Bostrom, 2014) suggests that advanced AI systems will pursue subgoals (self-preservation, resource acquisition, goal preservation) that may conflict with human interests regardless of which actor deploys them. Winner and losers alike face potential extinction or permanent disempowerment.
Crucially, the tragedy persists because the rate of capability gains in a race condition naturally outpaces the rate of safety verification. Consequently, the winner is statistically likely to deploy an unsafe system before establishing control, rendering the expected utility of “winning” negative even for the victor. This conclusion is reinforced by empirical data: Bengio et al. (2025); Grace et al. (2018) found that substantial minorities of AI researchers estimate non-negligible probabilities of extremely bad outcomes including human extinction. If these probability estimates are even roughly accurate, the astronomical loss associated with an unsafe deployment outweighs even enormous competitive gains, making the race irrational for all participants.
Second, winner identity is uncertain ex ante. While racing, no actor knows with certainty they will be the winner. Multiple actors have credible capabilities. Geopolitical and corporate competition involves numerous sophisticated actors, any of whom might achieve breakthroughs. The “DeepSeek Shock” of early 2025 serves as a potent empirical validator of this uncertainty. By achieving frontier-level reasoning capabilities with orders-of-magnitude greater training efficiency, a resource-constrained actor demonstrated that massive capital and hardware “moats” are not impregnable (DeepSeek-AI, 2025). This reintroduced volatility into the race, proving that “winner” status is fragile and susceptible to algorithmic disruption. From behind the veil of this uncertainty, all actors face expected losses from racing even if the ex post winner (whoever that turns out to be) might retrospectively have benefited.
@@
Third, even decisive technological leads may be temporary. Unlike traditional military conquest where territory captured remains under control, AI capabilities can potentially be copied, stolen, or independently rediscovered. Winner-take-all assumptions may prove mistaken if multiple actors achieve advanced AI within relatively short timeframes, or if AI capabilities prove impossible to monopolize due to espionage, independent discovery, or proliferation.
@@
Critical caveat: This defense of Pareto-inefficiency relies on two empirical claims: (1) catastrophic risks from advanced AI are sufficiently probable and severe to dominate even winner-take-all competitive gains, and (2) actors incorporate these risks into their decision-making rather than pure output or power maximization. The first claim is defended by substantial technical literature on AI safety (Bostrom, 2014; Russell, 2019; Amodei et al., 2016). The second claim is more tenuous—if actors systematically discount existential risks or believe they can control advanced AI despite expert warnings, the tragedy structure may indeed break for prospective winners, transforming the coordination problem into a more traditional conquest dynamic.
@@ 5.1.5 Condition 5: Enforcement Difficulty
Monitoring Infeasibility: Even with agreements, monitoring compliance faces severe challenges. What constitutes “dangerous” AI capabilities may be ambiguous until after development. The distributed nature of AI research across many institutions resists centralized monitoring.
@@ 5.1.6 Result:
This analysis assumes actors incorporate existential risks into rational calculations, consistent with survey evidence showing substantial minorities of AI researchers estimate non-trivial extinction probabilities (Bengio et al., 2025; Grace et al., 2018). Actors who systematically discount such risks would not face the structural tragedy we identify, transforming the problem from coordination failure to conquest dynamics. However, the empirical evidence suggests leading actors do recognize these risks even while competitive pressure compels continued development—precisely the tragedy structure our framework predicts.
Robustness of the Racing Dynamic.
Just as with productivity, this tragedy structure is robust to relaxed assumptions. The dominance of acceleration holds in continuous strategy spaces: if actors choose a level of investment or compute $k_{i}\in[0,K_{max}]$, the marginal strategic advantage of increasing $k_{i}$ remains positive as long as competitors are advancing. Consequently, “slowing down” is not a stable equilibrium; the Nash equilibrium collapses to maximum feasible acceleration ($K_{max}$). Furthermore, strategic risk aversion (fear of falling behind) amplifies the race. Because the loss of sovereignty or market relevance is viewed as a certain, immediate catastrophe, while safety failure is viewed as a probabilistic, future catastrophe, risk-averse actors rationally prioritize avoiding the immediate deterministic loss, thereby intensifying the acceleration.
@@
Table 3 summarizes the comparison across eight crucial dimensions.
Table 3: AI Governance vs. Historical Arms Control Across Eight Dimensions
Dimension Chemical / Biological Nuclear AI
@@ 5.2.8 Dimension 8: Economic-Military Convergence
Chemical and biological weapons converged only military considerations. Nations developed and deployed them (when they did) for military advantage, making the coordination problem purely security-focused.
Nuclear weapons similarly converged only military considerations, with the exception of some civilian nuclear power development which proved separable.
@@ 5.3 Empirical Validation: The Russia-Ukraine Drone War
The Russia-Ukraine conflict validates the framework’s structural predictions, not for autonomous AI specifically, but for the competitive dynamics that will govern its development. When Russia invaded Ukraine in February 2022, both sides had minimal drone capabilities. Within two years, deployment escalated from dozens to thousands of drones monthly (Center for Strategic and International Studies, 2024; Watling and Reynolds, 2023), with Ukraine expected to produce 4.5 million drones in 2025 (Nieczypor and Matuszak, 2025). This exponential growth occurred despite years of prior international dialogue on weapons governance. While current drones remain predominantly remotely piloted rather than AI-controlled, the success of drone warfare in this and other recent conflicts is expected to increase drone use in future warfare (Fukuyama, 2024), with autonomous AI-controlled systems a likely next stage as capabilities mature.
@@
The escalation pattern demonstrates our framework’s core dynamic: once a technology confers decisive military advantage under existential stakes, competitive pressure compels rapid adoption regardless of prior governance intentions. Each side faced overwhelming incentive to match the other’s deployment (C3: dominance driven by existential stakes). Drone deployment by one side created immediate pressure for counter-deployment (C2: negative externalities). No international authority could enforce restraint during active conflict (C5: enforcement impossibility). Pre-war governance discussions proved irrelevant once competitive pressure intensified—precisely the trajectory our framework predicts for autonomous AI systems as capabilities mature.
This case validates key structural predictions: when existential stakes are present, competitive dynamics overwhelm governance aspirations; escalation occurs rapidly when technology enables it; dual-use technology resists governance attempts; and verification and enforcement fail during intense competition. The drone war thus serves as both confirmation of the framework and a leading indicator for AI governance challenges ahead.
5.4 Why Standard Solutions Face Structural Barriers
Our framework explains why proposed AI governance solutions face fundamental challenges:
Solution 1: Voluntary Pauses
Problem: Violates dominance condition. As noted in the context of military AI arms control, strategic stability concerns make voluntary restraint inherently unstable (Maas, 2019). Consequently, the 2023 open letter calling for an AI pause produced no actual pause despite thousands of signatures (Future of Life Institute, 2023).
Solution 2: International Treaties
Problem: Violates enforcement condition. Treaties face international anarchy plus verification impossibility given software’s intangible nature. Unlike nuclear non-proliferation (which had high entry barriers and decades for institution-building), AI lacks favorable conditions.
Solution 3: Laboratory Self-Governance
Problem: Violates dominance condition. Safety-conscious labs face first-mover disadvantage as they may be eliminated by competitive pressure, creating tragic selection for less cautious actors.
Solution 4: National Regulation
Problem: Violates enforcement condition. AI development can relocate to permissive jurisdictions. No single nation can regulate global development. Race-to-bottom dynamics undermine unilateral regulation.
Solution 5: Technical Safety Research
Problem: While crucial, technical safety research cannot solve the coordination problem. Even with technically safe systems, competitive pressure drives deployment before adequate verification. Safety becomes a constraint that competitive actors work around rather than a prerequisite they satisfy.
Solution 6: Compute Governance
Problem: Restricting access to specialized AI hardware represents the most physically verifiable intervention point, as advanced chips provide a tangible chokepoint that software development lacks. However, this advantage erodes under competitive pressure: nations race to develop domestic semiconductor capacity, algorithmic efficiency reduces hardware requirements over time, and the dual-use nature of computing infrastructure prevents meaningful restrictions without crippling broader economic development. Compute governance may offer a temporary window but faces the same structural dynamics—competitive pressure overwhelming coordination—that the framework identifies across all five conditions.
Solution 6: Energy Monitoring
@@ 5.4 Why Standard Solutions Face Structural Barriers
Problem: Energy monitoring faces similar dynamics. Data center energy consumption is observable and growing rapidly, suggesting energy infrastructure as a potential verification mechanism. However, nations must invest in energy infrastructure for reasons entirely independent of AI such as grid modernization, electrification, industrial policy. Restricting energy investment to limit AI development is economically and politically infeasible. Any agreement to monitor energy for AI governance purposes would be trivially circumvented by dual-use justifications.
These solutions fail because they leave underlying conditions intact. Only interventions that break at least one of the five conditions have prospects for success.
5.5 What Would Work
Based on our framework, effective interventions must break at least one of the five conditions. For AI governance, this is structurally more difficult than for productivity due to amplified intensity across all conditions.
Break Condition 1 (N-Player): Consolidate AI development in a single international consortium or heavily regulated monopoly. This faces severe feasibility constraints: national security concerns prevent international consolidation, while domestic monopoly risks concentrating existential-risk technologies in single entities. The structural barriers are formidable.
Break Condition 2 (Externalities): Eliminate the externality where one actor’s AI advancement threatens others. This appears structurally impossible since AI capabilities inherently create strategic advantages that disadvantage competitors, whether in economic or military domains.

### v3

@@ 5 The Tragedy of AI Governance
Scope of Analysis:
Our analysis focuses on the governance challenge posed by transformative AI breakthroughs: systems approaching or exceeding human-level general intelligence (AGI) or superintelligence. We assume such systems pose non-excludable existential risks if developed without adequate safety measures (Bostrom 2014, Russell 2019, Bengio 2025). We examine the implications of relaxing this assumption after establishing Pareto-inefficiency (C4).
@@ 5.1.3 Condition 3: Dominance Property
The dominance is driven by existential stakes for firms, nations, and researchers. For example, a firm who cannot raise capital for massive compute (or acceleration) risk obsolescence. This structural pressure forced even the leading non-profit actor to restructure first into a ‘capped-profit’ entity (OpenAI, 2019) and eventually into a Public Benefit Corporation to remove fundraising limits (OpenAI, 2025).
@@ 5.1.4 Condition 4: Pareto-Inefficiency
Asymmetric Gains and the Winner-Take-All Problem:
AI development creates potential for extremely asymmetric outcomes. Unlike productivity competition where firms maintain roughly proportional positions, AI capabilities races may produce winner-take-all dynamics where the first to achieve advanced AI captures decisive advantages, potentially controlling entire markets, achieving strategic dominance, or even technological singularity.
@@
First, existential risk is non-excludable.
The catastrophic risks from misaligned or insufficiently robust advanced AI systems affect all actors regardless of who develops them first (Bostrom, 2014). A winner who achieves advanced AI through hasty development faces the same risks of loss of control, unintended consequences, or catastrophic failure as any other actor. The instrumental convergence thesis (Bostrom, 2014) suggests that advanced AI systems will pursue subgoals (self-preservation, resource acquisition, goal preservation) that may conflict with human interests regardless of which actor deploys them. Winners and losers alike face potential extinction or permanent disempowerment.
Crucially, the tragedy persists because the rate of capability gains in a race condition naturally outpaces the rate of safety verification. Consequently, the winner is statistically likely to deploy an unsafe system before establishing control, rendering the expected utility of ‘winning’ negative even for the victor. This conclusion is reinforced by empirical data: Bengio et al. (2025); Grace et al. (2018) found that substantial minorities of AI researchers estimate non-negligible probabilities of extremely bad outcomes including human extinction. If these probability estimates are even roughly accurate, the astronomical loss associated with an unsafe deployment outweighs even enormous competitive gains, making the race irrational for all participants.
Second, winner identity is uncertain ex ante.
While racing, no actor knows with certainty they will be the winner. Multiple actors have credible capabilities. Geopolitical and corporate competition involves numerous sophisticated actors, any of whom might achieve breakthroughs. The “DeepSeek Shock” of early 2025 serves as a potent empirical validator of this uncertainty. By achieving frontier-level reasoning capabilities with orders-of-magnitude greater training efficiency, a resource-constrained actor demonstrated that massive capital and hardware ‘moats’ are not impregnable (DeepSeek-AI, 2025). Crucially, while this lowers barriers to entry, it does not reduce the ‘winner-take-all’ prize of the final breakthrough; instead, it increases the volatility of the race by proving that current leaders cannot rely on resource dominance to secure their future position. The identity of the ultimate winner remains radically uncertain. From behind this veil of uncertainty, all actors—even current leaders—face expected losses from an uncoordinated race.
@@
Third, even decisive technological leads may be temporary.
Unlike traditional military conquest where territory captured remains under control, AI capabilities can potentially be copied, stolen, or independently rediscovered. Winner-take-all assumptions may prove mistaken if multiple actors achieve advanced AI within relatively short timeframes, or if AI capabilities prove impossible to monopolize due to espionage, independent discovery, or proliferation.
@@
Critical caveat:
This defense of Pareto-inefficiency relies on two empirical claims: (1) catastrophic risks from advanced AI are sufficiently probable and severe to dominate even winner-take-all competitive gains, and (2) actors incorporate these risks into their decision-making rather than pure output or power maximization. The first claim is defended by substantial technical literature on AI safety (Bostrom, 2014; Russell, 2019; Amodei et al., 2016). The second claim is more tenuous: if actors systematically discount existential risks or believe they can control advanced AI despite expert warnings, the existential dimension of the tragedy weakens.
However, even under systematic discounting of existential risk, the tragedy structure persists at reduced intensity. Coordination failure would manifest through wasteful resource competition: duplicated R&D spending across labs, unsustainable talent bidding wars, massive capital obligations against uncertain returns, and infrastructure investments vulnerable to algorithmic breakthroughs that devalue hardware moats-as DeepSeek’s efficiency gains demonstrated. The resulting “war of attrition” remains Pareto-inefficient (all participants would prefer coordinated investment), lowering C3 and C4 from extreme to high intensity. The enforcement difficulty (C5) remains at extreme intensity regardless of risk estimates, as the technical barriers to verification are unchanged. The tragedy persists; only its stakes diminish.
@@ 5.1.5 Condition 5: Enforcement Difficulty
Monitoring Infeasibility: Even with agreements, monitoring compliance faces severe challenges. What constitutes ‘dangerous’ AI capabilities may be ambiguous until after development. The distributed nature of AI research across many institutions resists centralized monitoring.
@@ 5.1.6 Result: Structural Tragedy Confirmed
As noted in the critical caveat above, actors who systematically discount existential risks would still face a structural tragedy at reduced intensity (Index $\approx 1,000$ rather than $\approx 3,981$), comparable to climate change and bank runs. The empirical evidence suggests leading actors do recognize catastrophic risks even while competitive pressure compels continued development, precisely the tragedy structure our framework predicts.
Robustness of the Racing Dynamic:
Just as with productivity, this tragedy structure is robust to relaxed assumptions. The dominance of acceleration holds in continuous strategy spaces: if actors choose a level of investment or compute $k_{i}\in[0,K_{max}]$, the marginal strategic advantage of increasing $k_{i}$ remains positive as long as competitors are advancing. Consequently, ‘slowing down’ is not a stable equilibrium; the Nash equilibrium collapses to maximum feasible acceleration ($K_{max}$). Furthermore, strategic risk aversion (fear of falling behind) amplifies the race. Because the loss of sovereignty or market relevance is viewed as a certain, immediate catastrophe, while coordination failure consequences (whether existential or economic) are viewed as probabilistic and future catastrophe, risk-averse actors rationally prioritize avoiding the immediate deterministic loss, thereby intensifying the acceleration.
@@
Table 3 summarizes the comparison across eight crucial dimensions. We now discuss each dimension in detail.
Table 3: AI Governance vs. Historical Arms Control Across Eight Dimensions. AI represents a structural outlier, facing the most unfavorable condition for coordination across all eight dimensions simultaneously, whereas historical cases benefitted from mitigating factors (e.g., observability, separability) in key areas.
Dimension Chemical / Bio. Nuclear AI
@@ 5.2.8 Dimension 8: Economic-Military Convergence
Chemical and biological weapons involved only military considerations. Nations developed and deployed them (when they did) for military advantage, making the coordination problem purely security-focused.
Nuclear weapons similarly involved only military considerations, with the exception of some civilian nuclear power development which proved separable.
@@ 5.3 Empirical Validation: The Russia-Ukraine Drone War
The Russia-Ukraine conflict validates the framework’s structural predictions for competitive dynamics under existential stakes. When Russia invaded Ukraine in February 2022, both sides had minimal drone capabilities. Within two years, deployment escalated from dozens to thousands of drones monthly (Center for Strategic and International Studies, 2024; Watling and Reynolds, 2023), with Ukraine expected to produce 4.5 million drones in 2025 (Nieczypor and Matuszak, 2025). This exponential growth occurred despite years of prior international dialogue on weapons governance.
@@
The escalation demonstrates our framework’s core dynamic. Each side faced overwhelming incentive to match the other’s deployment (C3: dominance driven by existential stakes). Drone deployment by one side created immediate pressure for counter-deployment (C2: negative externalities). No international authority could enforce restraint during active conflict (C5: enforcement impossibility). Pre-war governance discussions proved irrelevant once competitive pressure intensified.
Current drones remain predominantly remotely piloted: physically observable, with detectable manufacturing and verifiable deployment. Yet governance still failed. Now consider autonomous drones: more lethal, requiring fewer operators, capable of faster response, but also more prone to errors without human oversight. The governance challenge steepens: verifying autonomous compliance requires inspecting software logic, yet strategic incentives preclude sharing targeting code that would reveal vulnerabilities to adversaries. The transition from remote-controlled to autonomous systems amplifies every dimension of the coordination failure while adding verification impossibility.
If governance fails for remotely piloted drones, a relatively governable technology, it will fail more decisively for autonomous drones. And if autonomous drone governance proves intractable, consider AI integrated across the full spectrum of military systems: autonomous vehicles, cyber operations, logistics, intelligence analysis, command and control. Each integration multiplies the coordination challenge our framework identifies. The drone war thus serves as both confirmation of the framework and a conservative lower bound for AI governance challenges ahead.
5.4 Why Standard Solutions Face Structural Barriers?
Our framework explains why commonly proposed AI governance solutions face structural barriers. Each solution fails to adequately address one or more of the five necessary conditions.
Solution 1 — Voluntary Industry Self-Regulation:
Encourage AI labs to voluntarily adopt safety standards and slow development.
Problem: Violates dominance condition (C3). As Maas (2019) argues in the context of military AI arms control, strategic stability concerns make voluntary restraint inherently unstable—any actor that restrains development while competitors advance faces strategic elimination. The 2023 open letter calling for an AI development pause produced no actual pause despite thousands of signatures (Future of Life Institute, 2023), demonstrating that even widespread recognition of risks cannot overcome competitive pressure. Even safety-conscious organizations face overwhelming pressure to accelerate when rivals make breakthroughs. OpenAI’s trajectory exemplifies this: founded as a non-profit to ensure safe AI development, competitive necessity forced restructuring first into a “capped-profit” entity (OpenAI, 2019) and eventually into a Public Benefit Corporation to remove fundraising limits (OpenAI, 2025). Voluntary restraint is unstable under winner-take-all competitive dynamics.
Solution 2 — International Treaties:
Establish binding international agreements limiting AI capabilities development, analogous to nuclear non-proliferation treaties.
Problem: Violates enforcement condition (C5). As our eight-dimension comparison (Table 3) demonstrates, AI faces more severe verification challenges than biological or chemical weapons. Koblentz (2009) documents how verification failures have rendered the Biological Weapons Convention largely ineffective despite physical detectability of biological weapons programs. AI faces even steeper challenges: software can be developed covertly, training can be concealed, and capabilities cannot be reliably detected without access to proprietary systems. International anarchy means no global authority can enforce compliance, and verification impossibility precludes credible monitoring.
Solution 3 — AI Safety Research:
Solve the technical alignment problem, making advanced AI systems safe by design.
Problem: Insufficient to resolve coordination failure even if technically successful. Safety research addresses the risk of misaligned AI but cannot eliminate competitive pressure from aligned AI. As discussed in Condition 4 in Section 5.1, even perfectly aligned AI systems create winner-take-all dynamics through economic and strategic advantages. The coordination failure arises from competitive pressure to be first, which safety research alone cannot eliminate. Moreover, safety research itself faces the racing dynamic: labs may cut corners on safety validation when competitors approach breakthroughs (Armstrong et al., 2016).
Solution 4 — National Regulation:
Implement comprehensive safety requirements and capability restrictions within individual nations.
Problem: Violates enforcement condition (C5) through regulatory arbitrage. As Armstrong et al. (2016) formally model, unilateral safety regulations create competitive disadvantage, causing development to migrate to less cautious jurisdictions—merely shifting rather than reducing existential risk. The decentralized, software-based nature of AI development enables rapid relocation compared to physical industries. Without global coordination, national regulation faces the same race-to-bottom dynamics as labor standards under globalization.
Solution 5 — Compute Governance:
Restrict access to advanced computing hardware required for frontier AI development, as proposed by Anderljung et al. (2023).
Problem: Faces erosion over time despite initial promise. While compute restrictions could theoretically create entry barriers through hardware chokepoints, Sastry et al. (2024) document how algorithmic efficiency improvements systematically lower hardware requirements, enabling actors to achieve frontier capabilities with previously sub-frontier hardware. The DeepSeek breakthrough of early 2025 exemplifies this dynamic: achieving frontier reasoning capabilities with orders-of-magnitude greater training efficiency effectively undermines static compute-based thresholds. Over time, what requires cutting-edge datacenters today becomes achievable on commodity hardware, eroding the enforcement mechanism. Additionally, compute governance requires international cooperation that faces standard enforcement barriers from sovereignty concerns and verification challenges.
Solution 6 — Energy Monitoring:
Monitor data center energy consumption as a verification mechanism for frontier AI development, since large-scale training requires massive electricity usage. While Sastry et al. (2024) note that energy consumption correlates with compute usage for large training runs, they acknowledge this creates at best an indirect signal of AI development.
@@ 5.4 Why Standard Solutions Face Structural Barriers?
Problem: Energy monitoring faces fundamental dual-use problems. Data center energy consumption is observable and growing rapidly, but nations must invest in energy infrastructure for reasons entirely independent of AI—grid modernization, electrification, industrial policy, and general cloud computing demand. Restricting energy investment to limit AI development is economically and politically infeasible given these legitimate alternative uses. Any agreement to monitor energy for AI governance purposes would be trivially circumvented by dual-use justifications: large energy consumption could be attributed to cryptocurrency mining, scientific computing, commercial cloud services, or other legitimate activities. Unlike specialized AI chips that have limited alternative uses, energy infrastructure is inherently general-purpose, making verification through energy monitoring structurally impossible.
These solutions face structural barriers because they fail to address the complete set of conditions simultaneously. The analysis suggests that AI governance requires interventions of unprecedented scope and coordination—qualitatively beyond what successful arms control achieved—precisely because AI development satisfies all five tragedy conditions with extreme intensity.
5.5 What Would Work?
Our framework identifies interventions that could theoretically resolve the AI governance tragedy by breaking one or more of the five necessary conditions. For AI governance, this is structurally more difficult than for productivity due to amplified intensity across all conditions. We also emphasize that these interventions represent structural requirements, not policy recommendations, to clarify the magnitude of barriers any solution must overcome.
Break Condition 1 (N-Player Structure):
Consolidate all transformative AI development into a single global project—a heavily regulated monopoly or international consortium. Bostrom (2014) terms this a “Singleton”: a global coordination mechanism with sufficient authority to prevent competitive AI development. Such an entity could eliminate the racing dynamic by becoming the sole developer.
However, this solution creates severe risks that may exceed those it solves. Concentrating transformative AI development in a single entity enables permanent power consolidation, potential totalitarian control over transformative technology, and elimination of competitive safeguards against misuse. The cure might be worse than the disease, trading coordination tragedy for concentration risk. Additionally, achieving global agreement to establish such an entity faces formidable obstacles given competing national interests, sovereignty concerns, and deep mistrust among major powers regarding who would control such an entity.
Break Condition 2 (Binary Choice with Externalities):
Develop AI systems that inherently do not create strategic advantages, eliminating competitive pressure.
However, this is structurally difficult because AI capabilities inherently create strategic advantages since economic productivity, military capability, or scientific advancement cannot be developed in ways that do not affect competitive positions. This reflects the Collingridge (1980) dilemma: the strategic implications and competitive externalities of transformative AI only become fully apparent once the technology is developed, at which point competitive dependencies preclude coordinated restraint. Early restraint fails from uncertainty about implications; late restraint fails from path dependency and lock-in to competitive dynamics. The externality is intrinsic to the technology’s transformative potential.
Break Condition 3 (Dominance Property):